# LeapFrog: The Rowhammer Instruction Skip Attack

### v1

@@ Abstract
Since its inception, Rowhammer exploits have rapidly evolved into increasingly sophisticated threats not only compromising data integrity but also the control flow integrity of victim processes. Nevertheless, it remains a challenge for an attacker to identify vulnerable targets (i.e., Rowhammer gadgets), understand the outcome of the attempted fault, and formulate an attack that yields useful results.
In this paper, we present a new type of Rowhammer gadget, called a LeapFrog gadget, which, when present in the victim code, allows an adversary to subvert code execution to bypass a critical piece of code (e.g., authentication check logic, encryption rounds, padding in security protocols). The LeapFrog gadget manifests when the victim code stores the Program Counter (PC) value in the user or kernel stack (e.g., a return address during a function call) which, when tampered with, re-positions the return address to a location that bypasses a security-critical code pattern.
This research also presents a systematic process to identify LeapFrog gadgets. This methodology enables the automated detection of susceptible targets and the determination of optimal attack parameters. We first showcase this new attack vector through a practical demonstration on a TLS handshake client/server scenario, successfully inducing an instruction skip in a client application. We then demonstrate the attack on real-world code found in the wild, implementing an attack on OpenSSL. Our findings extend the impact of Rowhammer attacks on control flow and contribute to the development of more robust defenses against these increasingly sophisticated threats.
@@ 1 Introduction
The miniaturization of DRAM technology has inadvertently increased the susceptibility to bit flips and reliability issues. To mitigate data corruption, DRAM rows are refreshed at regular intervals, typically every 64 milliseconds. However, Kim et al. [25] discovered that rapid and repeated access to adjacent rows could accelerate charge leakage, leading to bit flips prior to the scheduled refresh, a phenomenon known as the Rowhammer effect [25]. Expanding on this, Seaborn et al. [43] demonstrated an even more efficient method known as double-sided Rowhammer, which exacerbates the issue.
@@
Further developments in exploiting the Rowhammer vulnerability have been numerous. Gruss et al. [19] achieved root privileges by flipping opcodes in the sudo binary through a single-location hammering technique. Additionally, the feasibility of launching Rowhammer attacks remotely via JavaScript was shown by Gruss et al. [20] and Ridder et al. [17]. Tatar et al. [45] and Lip et al. [31] extended the reach of Rowhammer to network-based attacks. Its applicability in cloud environments [51, 13] and on FPGA-CPU hybrid platforms [50] has also been demonstrated. Importantly, Kwong et al. [29] revealed that Rowhammer poses not only an integrity threat but also compromises confidentiality.
@@
Efforts to detect  [24, 12, 55, 23, 37, 21, 6, 16] and neutralize [20, 47, 8] Rowhammer attacks have been substantial. Yet, Gruss et al. [19] demonstrated the ineffectiveness of these countermeasures. Moreover, the security of ECC as a countermeasure was questioned by Cojocar et al. [14]. The Target Row Refresh (TRR) hardware countermeasure was recently circumvented as well, as shown by Frigo et al. [18] and further exploited by Ridder et al. [17] to target DDR4 chips with TRR. Notably, a recent study by Kogler et al. [27] highlighted the feasibility of hammering beyond adjacent locations to bypass TRR defenses.
@@
Software exploits enabled by Rowhammer were studied further in recent studies. Tobah et al. [1] introduced the notion of Rowhammer gadgets and a specialized attack. Specifically, if a victim code is designed to return benign data to an unprivileged user, and uses nested pointer dereferences, Rowhammer can be used to flip these pointers thereby gaining arbitrary read access in the victim’s address space. Adiletta et al. [3] demonstrated that even CPU internal elements like register values, which are occasionally saved to the stack, can be vulnerable to Rowhammer when they are temporarily stored in the stack and flushed to memory. Upon reloading, these corrupted values are returned to the registers, potentially leading to execution of faulty stack variables and security breaches.
@@
The threat of physical fault injection attacks has been acknowledged in the cryptographic community for some time [7]. For instance, OpenSSL incorporated error checks in CRT-based exponentiation early on to combat Bellcore attacks [7]. However, fault injection techniques have successfully compromised Elliptic Curve Parameters in the OpenSSL library [44]. Similarly, a Rowhammer-induced fault in WolfSSL, leading to ECDSA key exposure, was revealed in [35]. The vulnerability occurred during the TLS handshake process, involving the signing operation with private ECC keys. WolfSSL responded by introducing WOLF_SSL_CHECK_SIG_FAULTS, a series of checks during the signing stages to detect data tampering [36].
@@
A recent work by Adiletta et al. [3] targeted sensitive stack variables via Rowhammer threatening data integrity. In this paper we instead target the control-flow-integrity (CFI) and subvert the execution flow for malicious ends, e.g. to bypass sensitive sections of code user authentication and data encryption. For this we introduce LeapFrog a new Rowhammer attack vector that targets the PC when stored in the stack during function calls and context switches. Not all PC manipulations will yield useful results, as some jumps within the code will result in errors, like segfaults, or simply will not bypass the intended code logic. To explore the massive attack surface, we introduce an automated tool that dynamically analyzes code to detect this particular type of Rowhammer gadget [1].
@@ 2 Background
Rowhammer DRAM is stored in an array architecture of memory cells, where a capacitor and a transistor form each cell in the array. Each cell, capable of storing a value of either 1 or 0, is connected along word lines that extend across the row. Additionally, bit lines intersect the word lines perpendicularly, linking them to each cell. When the bit lines are brought into opposition (where one goes high and one goes low), positive feedback from a sense amplifier sets the state of the cell to be high or low. The sense amplifier consists of two cross connected inverters between the cells. For the cell to retain its state, the sense amplifier must be disconnected [2].
@@
The Rowhammer attack works by abusing the sense amplifiers ability to set the state of the cell. The capacitor will leak voltage, and must undergo a refresh of voltage every 64 ms (or less) according to standard JETEC convention. Reading the cell introduces noise into the system due to fluctuating voltages, and the noise can be amplified by the sense amp. As semiconductor technology improves, transistors shrink in size and the operating voltage also shrinks. The resulting cells are then able to operate at a higher speed, and the noise margins are reduced. Shrinking in size increases the signal coupling across traces and devices and magnifies crosstalk. When combined with the lower operating voltage and sharper edges due to higher speeds, the ratio of the crosstalk to the supply voltage increases significantly as technology improves making it easier for an adversary to exploit this type of attack. These can result in errors being written to the DRAM [2].
Instruction Skipping Instruction-skipping attacks are a type of fault attack that targets the normal execution flow of a program, particularly in embedded systems and secure circuits. Historically these techniques often needed physical access due to the timing and precision required to skip instructions. For example, laser fault injection has been demonstrated in inducing instruction skips in AES (Advanced Encryption Standard) that result in leaking private encryption keys [10]. In another example [41], researchers demonstrate how electromagnetic fault injection can effectively induce instruction skipping in the ARMv7-M architecture, specifically targeting AES.
@@
On the other hand, the kernel stack operates in kernel mode, a privileged mode of operation for the system’s kernel. Each thread of a process has its own kernel stack. This stack is used when the process executes system calls or when it is interrupted and the kernel needs to perform operations on behalf of the process. The kernel stack handles lower-level, system-critical operations like interrupt handling, system call implementation, and managing hardware interactions. It is kept separate from the user stack for security and stability, ensuring that user processes cannot directly access or interfere with kernel-level operations.
Data stored in the kernel stack includes CPU context for system calls, interrupt state information, and other kernel-specific data, while the user stack holds user-level process data like function calls and local variables. This separation reinforces the security and stability of the operating system by isolating user applications from the core kernel functions.
@@
Program Counters Program Counters (PCs), also known as Instruction Pointers, hold the memory address of the next instruction to be executed by the CPU. This mechanism ensures that instructions are executed in the correct sequence. The value of the Program Counter is typically not stored in the stack; rather, it’s stored in a dedicated register within the CPU. During the execution of a program, the PC is automatically incremented after each instruction is fetched, pointing to the subsequent instruction. However, during certain operations like function calls and interrupts, the PC value may be changed abruptly to a new address. In such cases, the return address (the original PC value) is often stored in the stack to enable the program to return to the correct point in the program after the operation is complete. For user function calls, the PC value gets pushed to the user stack, but if there is an exception, signal handler, or system call, the PC gets pushed to the kernel stack. This mechanism facilitates the smooth flow of program execution.
@@
Process Degradation Process degradation in computing refers to the intentional slowing down of a processor to create favorable conditions for certain types of attacks. A notable contribution in this field, HyperDegrade [4], combines previous approaches [5] with the use of simultaneous multithreading (SMT) architectures to significantly slow down processor performance, achieving a slowdown that is orders of magnitude greater than previous methods. It utilizes collateral Self Modifying Code (SMC) events to induce "machine clears", where the entire CPU pipeline is flushed, resulting in severe performance penalties. This process is triggered by cache line eviction, causing the invalidation of instructions in the victim’s L1 instruction cache, which the CPU may interpret as an SMC event. This mechanism amplifies the degradation effect, as instructions are sometimes fetched multiple times, leading to substantial slowdowns in CPU performance. This slowdown enhances the time granularity for FLUSH+RELOAD [22] attacks, enabling more effective exploitation of side-channel vulnerabilities in systems. The attack not only explores the implementation of this technique but also investigates the root causes of performance degradation, particularly focusing on cache eviction. Their findings have substantial implications in the realm of cryptography, as evidenced by the amplification of the Raccoon attack [33] on TLS-DH key exchanges and other protocols.
@@ 3 Related Work
Similar work [19] achieved privilege escalation through opcode flipping. Researchers loaded the sudo binary into memory from user space and flipped a bit in the opcode of the binary such that an incorrect password would cause authentication and the correct password would cause mis-authentication. They mention flipping conditional jumps that change program execution flow. While our work also can result in privilege escalation, we do so by attacking the process during run time and forcing the process to jump to an unexpected line of code by flipping the PC value.
To attack the binary during runtime, we had to overcome timing challenges as well as different detection problems to find vulnerable areas in the code. In [19] researchers used mmap to map the target binary into a vulnerable page in memory, demonstrating how memory waylaying and memory chasing techniques can force the mapped binary into the target page. This attack can potentially be mitigated by making the process execute only, and thus cannot be mapped with the mmap command. In contrast, our work can attack binaries that are unreadable from userland and are execute only. Additionally, our attack works on fundamentally different mechanics, so targets not susceptible to [19] may be susceptible to ours.
Another related work [1] demonstrates that code using nested pointer dereferences can corrupt bits in these pointers to reveal data to an unprivileged user. They demonstrate this vulnerability on ioctl given they can flood the kernel heap with data by spawning processes (a method they call "spraying"), increasing the probability a single bit-flip will point to malicious data in the heap that points to the location of secret data. Our work compliments and improves upon this prior work by increasing the number of vulnerable code patterns, since their work relies on the presence of specific code patterns that may not be present in the victim code.
Lastly, [54] demonstrates a Rowhammer attack methodology where researchers emulated Rowhammer bitflips on targets. They introduced the idea of simulating a flip in the EIP register value in the stack, which can force the execution to jump from kernel code to user code, like the ret2usr attack [28]. However, attacks that cause privilege escalation by jumping from kernel code to user code are mitigated by SMAP [15], which prevents the kernel from executing userland instructions. Our attack forces a process to jump within its own code space and privilege space, and thus is not affected by SMAP and introduces attack surfaces on new code patterns.
@@ 4 Threat Model
Similar to other Rowhammer attacks we assume the attacker is co-located on the same system as the victim  [25, 19, 51, 13, 20]. Co-location is a common threat model for many micro-architectural side-channel attacks and fault attacks  [30, 26, 11, 46, 48]. We do not assume root privilege or physical access to the machine. For some attacks, we assume that the attacker can send signals to pause it during the attack, such as sudo. However, we do not need to be root to send signals. Additionally, we assume that the system has TRR enabled, and bypass TRR with a multi-sided attack [18].
@@ 5 LeapFrog Attack
Figure 1: LeapFrog gadget in TLS handshake $addr_{src}$, the PC value that fault is injected into, is highlighted in blue. The new value after the fault injected, is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
Central to our investigation is the concept of LeapFrog gadgets, an important element in the orchestration of Rowhammer-based attacks targeting the Program Counter (PC). These gadgets are exploitable in scenarios where a process undergoes a context switch or executes a function call, leading to the storage of the PC value in either the kernel or user stack. The ingenuity of LeapFrog gadgets lies in their susceptibility to Rowhammer-induced bit flips due to them being stored in DRAM, enabling an attacker to subtly alter the PC value. This manipulation is designed to redirect the execution flow to a different code segment, ideally with minimal bit changes due to the blunt nature of Rowhammer and the higher probability of finding a faulty memory location with few or one faulty bits in the right location. In this paper, we assume that we can successfully find 1 bit flip within a page that is in the right location to fault the PC value to force the intended instruction skip.
In Figure 1 the storage of the PC value occurs in the kernel stack during the execution of wait_receive. In this scenario, a malicious server can hold the client process at the wait_receive function while hammering the PC value to force the process to jump to a new location upon returning from the function. In our assembly code analysis in Figure 1, we observe the original PC value is an address 0x555555555478. Through strategic bit flips, this value can be altered to 0x555555555578, effectively enabling an instruction skip (skipping one or more instructions) and jumping from the function call in wait_receive directly to a later point in the execution, bypassing the critical server authentication check. The practicability of such attacks, however, hinges on the feasibility of achieving the desired bit flips, a central challenge to the effectiveness of LeapFrog gadgets in real-world scenarios. In this scenario, flipping the PC from 0x555555555478 to 0x555555555578 only require 1 bit flip, which is a reasonble assumption for Rowhammer.
However, tiny variations in the C code can change the resulting assembly code significantly. For example, consider the first approach for a TLS handshake, where the process allocates memory for a message to be signed. The C code and its corresponding assembly code are shown in Listing 1.
@@
Alternatively, using a different method to allocate memory for the message results in a variation in the assembly code. This alternative approach and its corresponding assembly code are presented in Listing 2.
In the code space, the alternative approach (Listing 2) takes 0x55555555541f - 0x555555555413 or 12 bytes of instructions, while the original approach (Listing 1) occupies 8 bytes of assembly instruction. Given the assumption that only one bit per page can be reliably flipped, identifying useful instruction skips that require a single bit change, as illustrated in Figure 2, is crucial. This example illustrates the challenge in manually inspecting source code to determine the impact of tiny variations on assembly instruction distances. Hence, profiling binaries becomes an important tool in this context.
Figure 2: The best LeapFrog gadgets require a single bit flip, where the distance between the two lines of code is a power of 2

### v2

@@ Abstract
Since its inception, Rowhammer exploits have rapidly evolved into increasingly sophisticated threats compromising data integrity and the control flow integrity of victim processes. Nevertheless, it remains a challenge for an attacker to identify vulnerable targets (i.e., Rowhammer gadgets), understand the outcome of the attempted fault, and formulate an attack that yields useful results.
In this paper, we present a new type of Rowhammer gadget, called a LeapFrog gadget, which, when present in the victim code, allows an adversary to subvert code execution to bypass a critical piece of code (e.g., authentication check logic, encryption rounds, padding in security protocols). The LeapFrog gadget manifests when the victim code stores the Program Counter (PC) value in the user or kernel stack (e.g., a return address during a function call) which, when tampered with, repositions the return address to a location that bypasses a security-critical code pattern.
This research also presents a systematic process to identify LeapFrog gadgets. This methodology enables the automated detection of susceptible targets and the determination of optimal attack parameters. We first show the attack on a decision tree algorithm to show the potential implications. Secondly, we employ the attack on OpenSSL to bypass the encryption and reveal the plaintext. We then use our tools to scan the Open Quantum Safe library and report on the number of LeapFrog gadgets in the code. Lastly, we demonstrate this new attack vector through a practical demonstration in a client/server TLS handshake scenario, successfully inducing an instruction skip in a client application. Our findings extend the impact of Rowhammer attacks on control flow and contribute to developing more robust defenses against these increasingly sophisticated threats.
@@ 1 Introduction
The miniaturization of DRAM technology has inadvertently increased the susceptibility to bit flips and reliability issues. To mitigate data corruption, DRAM rows are refreshed at regular intervals, typically every 64 milliseconds. However, Kim et al. [30] discovered that rapid and repeated access to adjacent rows could accelerate charge leakage, leading to bit flips before the scheduled refresh, a phenomenon known as the Rowhammer effect [30]. Expanding on this, Seaborn et al. [53] demonstrated an even more efficient method known as the double-sided Rowhammer, which exacerbates the issue.
@@
Further developments in exploiting the Rowhammer vulnerability have been numerous. Gruss et al. [22] achieved root privileges by flipping opcodes in the sudo binary using a single-location hammering technique. Furthermore, Gruss et al. [23] and Ridder et al. [18] demonstrated the feasibility of launching Rowhammer attacks remotely via JavaScript. Tatar et al. [57] and Lip et al. [36] extended the reach of Rowhammer to network-based attacks. Its applicability has also been demonstrated in cloud environments [65, 14] and on hybrid FPGA-CPU platforms [64]. Importantly, Kwong et al. [34] revealed that Rowhammer poses not only an integrity threat but also compromises confidentiality.
@@
Efforts to detect  [27, 13, 69, 26, 45, 24, 5, 17] and neutralize [23, 61, 9] Rowhammer attacks have been substantial. However, Gruss et al. [22] demonstrated the ineffectiveness of these countermeasures. Moreover, Cojocar et al. [15] questioned the security of ECC as a countermeasure. The Target Row Refresh (TRR) hardware countermeasure was recently also circumvented, as shown by Frigo et al. [21] and further exploited by Ridder et al. [18] to target DDR4 chips with TRR. In particular, a recent study by Kogler et al. [32] highlighted the feasibility of hammering beyond adjacent locations to bypass TRR defenses.
@@
The software exploits enabled by Rowhammer were further studied in recent research. Tobah et al. [58] introduced the notion of Rowhammer gadgets and a specialized attack. Specifically, if a victim code is designed to return benign data to an unprivileged user, and uses nested pointer dereferences, Rowhammer can be used to flip these pointers thereby gaining arbitrary read access in the victim’s address space. Adiletta et al. [2] demonstrated that even internal CPU elements such as register values, which are occasionally saved to the stack, can be vulnerable to Rowhammer when they are temporarily stored in the stack and flushed to memory. Upon reloading, these corrupted values are returned to the registers, potentially leading to the execution of faulty stack variables and security breaches.
@@
The threat of physical fault injection attacks has been acknowledged in the cryptographic community for some time [8]. For instance, OpenSSL incorporated error checks in CRT-based exponentiation early on to combat Bellcore attacks [8]. However, fault injection techniques have successfully compromised Elliptic Curve Parameters in the OpenSSL library [56]. Similarly, Rowhammer-induced fault attacks in WolfSSL, leading to ECDSA key exposure, were revealed in [43, 19]. The vulnerabilities occurred during the TLS handshake process, involving the signing operation with private ECC keys. WolfSSL responded by introducing WOLF_SSL_CHECK_SIG_FAULTS and WOLFSSL_BLIND_PRIVATE_KEY, a series of checks during the signing stages to detect data tampering [40, 41].
@@
A recent work by Adiletta et al. [2] targeted sensitive stack variables via Rowhammer threatening data integrity. In this paper, we instead target the control-flow-integrity (CFI) and subvert the execution flow for malicious ends, e.g. to bypass sensitive sections of code user authentication and data encryption. For this, we introduce LeapFrog a new Rowhammer attack vector that targets the PC when stored in the stack during function calls and context switches. Not all PC manipulations will yield useful results, as some jumps within the code will result in errors, like segfaults, or simply will not bypass the intended code logic. To explore the massive attack surface, we introduce an automated tool that dynamically analyzes code to detect this type of Rowhammer gadget [58].
@@ 1.1 Our contributions
  3. 3.
    We scan the Open Quantum Safe library signature scheme, OpenSSL encryption, and a machine learning model - and quantify the number of potential LeapFrog gadgets present in the code.
  4. 4.
@@ 2 Background
Rowhammer DRAM is stored in an array architecture of memory cells, where a capacitor and a transistor form each cell in the array. Each cell, capable of storing a value of either 1 or 0, is connected along word lines that extend across the row. Additionally, bit lines intersect the word lines perpendicularly, linking them to each cell. When the bit lines are brought into opposition (where one goes high and one goes low), positive feedback from a sense amplifier sets the state of the cell to be high or low. The sense amplifier consists of two cross-connected inverters between the cells. For the cell to retain its state, the sense amplifier must be disconnected [1].
@@
The Rowhammer attack works by abusing the sense amplifier’s ability to set the state of the cell. The capacitor will leak voltage and must undergo a refresh of voltage every 64 ms (or less) according to the standard JETEC convention. Reading the cell introduces noise into the system due to fluctuating voltages, and the noise can be amplified by the sense amp. As semiconductor technology improves, transistors shrink in size and the operating voltage also shrinks. The resulting cells are then able to operate at a higher speed, and the noise margins are reduced. Shrinking in size increases the signal coupling across traces and devices and magnifies crosstalk. When combined with the lower operating voltage and sharper edges due to higher speeds, the ratio of the crosstalk to the supply voltage increases significantly as technology improves making it easier for an adversary to exploit this type of attack. These can result in errors being written to the DRAM [1].
Instruction Skipping Instruction-skipping attacks are a type of fault attack that targets the normal execution flow of a program, particularly in embedded systems and secure circuits. Historically these techniques often needed physical access due to the timing and precision required to skip instructions. For example, laser fault injection has been demonstrated in inducing instruction skips in AES (Advanced Encryption Standard) that resulted in leaking private encryption keys [11]. In another example [51], researchers demonstrate how electromagnetic fault injection can effectively induce instruction skipping in the ARMv7-M architecture, specifically targeting AES.
@@
On the other hand, the kernel stack operates in kernel mode, a privileged mode of operation for the system’s kernel. Each thread of a process has its own kernel stack. This stack is used when the process executes system calls or when it is interrupted and the kernel needs to perform operations on behalf of the process. The kernel stack handles system-critical lower-level operations such as interrupt handling, system call implementation, and managing hardware interactions. It is kept separate from the user stack for security and stability, ensuring that user processes cannot directly access or interfere with kernel-level operations. Data stored in the kernel stack includes CPU context for system calls, interrupt state information, and other kernel-specific data, while the user stack holds user-level process data like function calls and local variables. This separation reinforces the security and stability of the operating system by isolating user applications from the core kernel functions.
@@
Program Counters Program Counters (PCs), also known as Instruction Pointers, hold the memory address of the next instruction to be executed by the CPU. This mechanism ensures that instructions are executed in the correct sequence. The value of the Program Counter is typically not stored in the stack; rather, it’s stored in a dedicated register within the CPU. During the execution of a program, the PC is automatically incremented after each instruction is fetched, pointing to the subsequent instruction. However, during certain operations like function calls and interrupts, the PC value may be changed abruptly to a new address. In such cases, the return address (the original PC value) is often stored in the stack to enable the program to return to the correct point in the program after the operation is complete. For user function calls, the PC value is pushed to the user stack, but if there is an exception, signal handler, or system call, the PC gets pushed to the kernel stack. This mechanism facilitates the smooth flow of program execution.
@@
Process Degradation Process degradation in computing refers to the intentional slowing down of a processor to create favorable conditions for certain types of attacks. A notable contribution in this field, HyperDegrade [3], combines previous approaches [4] with the use of simultaneous multithreading (SMT) architectures to significantly slow down processor performance, achieving a slowdown that is orders of magnitude greater than previous methods. It utilizes collateral Self Modifying Code (SMC) events to induce “machine clears”, where the entire CPU pipeline is flushed, resulting in severe performance penalties. This process is triggered by cache line eviction, causing the invalidation of instructions in the victim’s L1 instruction cache, which the CPU may interpret as an SMC event. This mechanism amplifies the degradation effect, as instructions are sometimes fetched multiple times, leading to substantial slowdowns in CPU performance. This slowdown enhances the time granularity for FLUSH+RELOAD [25] attacks, enabling more effective exploitation of side-channel vulnerabilities in systems. The attack not only explores the implementation of this technique but also investigates the root causes of performance degradation, particularly focusing on cache eviction. Their findings have substantial implications in the realm of cryptography, as evidenced by the amplification of the Raccoon attack [39] on TLS-DH key exchanges and other protocols.
@@ 3 Related Work
To attack the binary during runtime, we had to overcome timing challenges as well as different detection problems to find vulnerable areas in the code. In [22] researchers used mmap to map the target binary into a vulnerable page in memory, demonstrating how memory waylaying and memory-chasing techniques can force the mapped binary into the target page. This attack can potentially be mitigated by making the process execute only, and thus cannot be mapped with the mmap command. In contrast, our work can attack binaries that are unreadable from userland and are executed only. Additionally, our attack works on fundamentally different mechanics, so targets not susceptible to [22] may be susceptible to ours.
Another related work [58] demonstrates that code using nested pointer dereferences can corrupt bits in these pointers to reveal data to an unprivileged user. They demonstrate this vulnerability on ioctl given they can flood the kernel heap with data by spawning processes (a method they call “spraying”), increasing the probability a single bit-flip will point to malicious data in the heap that points to the location of secret data. Our work compliments and improves upon this prior work by increasing the number of vulnerable code patterns since their work relies on the presence of specific code patterns that may not be present in the victim code.
Lastly, [68] demonstrates a Rowhammer attack methodology where researchers emulated Rowhammer bitflips on targets. They introduced the idea of simulating a flip in the EIP register value in the stack, which can force the execution to jump from kernel code to user code, like the ret2usr attack [33]. However, attacks that cause privilege escalation by jumping from kernel code to user code are mitigated by SMAP [16], which prevents the kernel from executing userland instructions. Our attack forces a process to jump within its own code space and privilege space and thus is not affected by SMAP and introduces attack surfaces on new code patterns.
@@ 4 Threat Model
Similar to other Rowhammer attacks we assume the attacker is co-located on the same system as the victim  [30, 22, 65, 14, 23]. Co-location is a common threat model for many micro-architectural side-channel attacks and fault attacks  [35, 31, 12, 60, 62]. We do not assume root privilege or physical access to the machine. We only assume that the system has TRR enabled, and bypasses TRR with a many-sided attack [21].
@@ 5 LeapFrog Attack
Figure 1: LeapFrog gadget in TLS handshake $addr_{src}$, the PC value that fault is injected into, is highlighted in blue. The new value is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
LeapFrog gadgets are exploitable in scenarios where a process undergoes a context switch or executes a function call, leading to the storage of the PC value in either the kernel or user stack. The ingenuity of LeapFrog gadgets lies in their susceptibility to Rowhammer-induced bit flips due to them being stored in DRAM, enabling an attacker to alter the PC value subtly. This manipulation is designed to redirect the execution flow to a different code segment, ideally with minimal bit changes due to the blunt nature of Rowhammer and the higher probability of finding a faulty memory location with few or one faulty bits in the right location. In this paper, we assume that we can successfully find a 1-bit flip within a page that is in the right location to fault the PC value to force the intended instruction to skip.
In Figure 1 the storage of the PC value occurs in the kernel stack during the execution of wait_receive. In this scenario, a malicious server can hold the client process at the wait_receive function while hammering the PC value to force the process to jump to a new location upon returning from the function. In our assembly code analysis in Figure 1, we observe the original PC value is an address 0x555555555478. Through strategic bit flips, this value can be altered to 0x555555555578, effectively enabling an instruction skip (skipping one or more instructions) and jumping from the function call in wait_receive directly to a later point in the execution, bypassing the critical server authentication check. The practicability of such attacks, however, hinges on the feasibility of achieving the desired bit flips, a central challenge to the effectiveness of LeapFrog gadgets in real-world scenarios. In this scenario, flipping the PC from 0x555555555478 to 0x555555555578 only requires a 1-bit flip, which is a reasonable assumption for Rowhammer.
However, tiny variations in the C code can change the resulting assembly code significantly. For example, consider the first approach for a TLS handshake, where the process allocates memory for a message to be signed. The C code and its corresponding assembly code are shown in Listing 1. Alternatively, using a different method to allocate memory for the message results in a variation in the assembly code. This alternative approach and its corresponding assembly code are presented in Listing 2.
@@
In the source code space, the alternative approach (Listing 2) takes 0x55555555541f - 0x555555555413 or 12 bytes of instructions, while the original approach (Listing 1) occupies 8 bytes of assembly instruction (this excludes the size of the last instruction). Given the assumption that only one bit per page can be reliably flipped, identifying useful instruction skips that require a single bit change, as illustrated in Figure 2, is crucial. This example illustrates the challenge of manually inspecting source code to determine the impact of tiny variations on assembly instruction distances. Hence, profiling binaries becomes an important tool in this context.
Figure 2: The best LeapFrog gadgets require a single-bit flip, where the distance between the two lines of code is a power of 2.